(A saturated solution of calcium hydroxide was titraed with 0.005 M solution and it took 13.50 acid to reach neutralization point a) Caluclate the number of moles of acid used ) How may moles of hydroxide was neutralized ? c) What is the molar solubility in molarityof Calcium hidroxide ? ) Calculate the Ksp value calcium hydroxide
1
Expert's answer
2020-11-14T13:57:57-0500

We assume the volume of calcium hydroxide is 25mL

Mole of acid = 0.005 ×13.50×10-3 = 6.75×10-5 mole

Mole of base= mole of acid used = 6.75 ×10-5 mole

Molar solubility = mole / volume = 6.75×10-5 / 25×10-3 = 2.7×10-3M

Ksp= [Ca]2+ [OH-]2 = 2.7×10-3 × (2.7× 10-3) 2

= 1.968×10-8 M3